OpenStudy (anonymous):

Your agency is paid on a traditional 15 percent media commission. You purchase $300,000 in network TV time to run an ad for your client. How much will your client pay for the TV time? How much will the network receive? How much will the agency be paid? The ad you ran cost $30,000 to produce. Your agency charges a 17.65 percent markup for production costs. How much will your client pay for the production? How much will your agency receive?

OpenStudy (anonymous):

well how much will the client pay first of all

OpenStudy (anonymous):

30 000*(1,1765)= 35 295 =a 300 000*1,15=345 000 =b a+b=first question a+b-costs=2nd.

OpenStudy (anonymous):

1=100% 0,15=15% and so on.